OpenStudy (anonymous):

Find the surface area of a sphere with diameter 17 cm. A. 907.92 cm2 B. 226.98 cm2 C. 289 cm2 D. 106.81 cm2

OpenStudy (anonymous):

226.98 cm2

OpenStudy (callisto):

surface area = 4πr^2 where r = diameter/2 Can you try it now?

OpenStudy (anonymous):

Surface area of sphere = 4 pi r^2 = 4 pi (17/2)^2 = 907.9^2

OpenStudy (king):

S.A=4pi r^2 =4*22/7*8.5*8.5 =907.46

OpenStudy (king):

sry 907.92